OpenStudy (anonymous):

5. How long will it take an Amtrak passenger train to catch up to a freight train if the Amtrak’s speed is 80 mph, the freight train’s speed is 60 mph, and the freight train had an hour and a half head start?

OpenStudy (anonymous):

d=rt ... amtrak=80x f.t=60(1.5+x) ... 80x=90+60x

OpenStudy (anonymous):

SO it will take the Amtrak train 4 1/2 hours to catch up ?

OpenStudy (anonymous):

ya!